# Free energy barrier and thermal-quantum behavior of sliding bilayer graphene

### v3

@@ III.2 Shear frequency
Figure 3: Bilayer graphene phonon dispersion. The inset shows a zoom in close to $\Gamma$ to better visualize the doubly degenerate shear modes S at $\Gamma$, and the breathing mode B. Due to the high speed velocity in graphene, the region of the shear mode in the BZ is small (a fraction of $\Gamma$M), and is overrepresented in small supercell calculations. The breathing mode, on the other hand, is flat, so nearby sampling points are not necessary to converge the free energy.
Figure 4: Shear frequency as a function of temperature using SSCHA (blue circles) and QHA (red crosses). There is no experimental data for the temperature dependence of the shear frequency as a function of temperature for bilayer graphene. Experimental values correspond to bulk (blue crosses) and folded 2+2 graphene (orange circles), which are corrected by the factor of a nearest neighbor modelHe et al. (2016) to compare to the bilayer case (and rigidly shifted by -5 cm-1 to facilitate the visual comparison). The agreement with the temperature dependence of the bulk values is good.
Regarding the experimental data, in Ref. Cong and Yu, 2014, the measurements are done in 2+2 folded graphene (4 layers). In order to compare with our 2-layer calculations, we multiply by a correction factor using a nearest-neighbor model for layersHe et al. (2016). In the case of bulk, such a factor is $\sqrt{2}$ (the force exerted by a layer below and above, proportional to the square of the frequency, is doubled when considering only nearest neighbor interactions), which we use to compare to the bulk measurements of Ref. Mishina et al., 2000. The temperature dependence of both SSCHA and QHA curves is similar to the experimental adjusted bulk values. Folded graphene values have a steeper dependence but are likely unsuitable for comparison, because layers fold with different rotational angles, while layers are aligned in standard graphite. The shear frequency has also been measured in other materials like bilayer NbSe2 in Ref. He et al., 2016, or bulk $h$-BN in Ref. Cuscó et al., 2019, which show a similar temperature dependence (adjusting the bulk value by a $\sqrt{2}$ factor) of roughly -0.5 cm-1/100 K.
@@ III.3 Thermal expansion
Figure 5: (a) Graphite out-of-plane coefficient of TE as a function of temperature using SSCHA and QHA. PIMD calculationsHerrero and Ramírez (2020) and experimental data Bailey and Yates (1970) are also included. PIMD overall agrees well, while QHA and SSCHA (with the potential of Ref. Wen et al., 2018) underestimates the TE. (b) Graphite in-plane coefficient of TE using SSCHA and QHA. QHA again underestimates the dependence at low temperatures, and also at high temperatures. On the other hand, PIMD gives a dependence that is surprisingly very far from experimental values. SSCHA gives excellent results both at low and high temperatures, considering also the disagreement within experimental data.
To further check the adequacy of our method (the use of SSCHA with the in-planeLibbi et al. (2020) and out-of-planeWen et al. (2018) interatomic potentials), we look at the TE in the previous calculations and also in bulk graphite, and compare to experiments. The TE coefficients are given by $\alpha_{\mathrm{a}}(T)=\tfrac{1}{a}\tfrac{da}{dT}$ and $\alpha_{\mathrm{c}}(T)=\tfrac{1}{c}\tfrac{dc}{dT}$, with $a$ and $c$ the in-plane and out-of-plane lattice parameters, respectively.
In Fig. 5, we show the TE coefficient calculated with SSCHA and QHA in bulk, compared with experimental dataBailey and Yates (1970); Marsden et al. (2018) and PIMD calculations Herrero and Ramírez (2020). SSCHA results underestimate the TE. PIMD results are overall very good, although they also give values somewhat lower than experimentsBailey and Yates (1970); Marsden et al. (2018).
In-plane SSCHA calculations give excellent results. QHA works well at low temperatures but departs from experiments at higher temperatures. We conducted a careful convergence of these results, reported in the SI, Sec. S3. PIMD calculations of Ref. Herrero and Ramírez, 2020, although they work very well out-of-plane, greatly overestimate in-plane TE, presumably due to the inadequacy of their interatomic potential. It would be interesting in future work to compare SSCHA and PIMD using the same potentials.
@@ Acknowledgments
Supplementary Information
S1 Barrier
To determine the barrier, the free energy has to be determined at the equilibrium position and SP. At the equilibrium position, there are two shear modes, indicated with $\mathbf{S}$ in Fig. 3(b). The mode perpendicular to the bond direction shown in Fig. 1 is stable in SP, while the shear mode along the reaction coordinate (using the language of chemical reactions) is unstable. As opposed to the breathing mode, that is flat around $\Gamma$, the shear modes are only present in a small region of the BZ ($1/N^{2}$, with $N$ larger than 20), as a result of the large speed of sound in graphene. So in small supercells, the shear modes will be over-represented. This is more problematic at SP, where the SSCHA frequency corresponding to she unstable mode takes very small values (see Fig. S1), which makes the convergence of the free energy hard, mostly due to the logarithm term in $F_{\mathcal{H}}$ in Eq. (4). Just determining the barrier with this equation and a $N\times N\times 1$ supercell, results in Fig. S2, which is not well converged. All SSCHA calculations used 24000 configurations, both for the barrier and the TE.
In order to get better convergence with $N$, we use the interpolation method mentioned in Sec. II.2, and obtain a much more stable barrier as a function of $1/N$ (see Fig. S3). The $40\times 40$ interpolation grid used for $F_{\mathcal{H}}$ corresponds to a supercell of 6400 atoms. For 100, 200 and 300 K, the barrier seems well converged. For a 1000 K, the trend seems to indicate convergence within 0.02 meV/atom, so for the intermediate temperatures 400 and 500 K, the barrier is likely also converged within a similar error, resulting in Fig. 2 of the main text.
To better understand the reduction in the barrier, we show in Fig. S4 the SSCHA frequencies (around the shear modes) at equilibrium and SP at $T=100$ K. The SSCHA frequencies at equilibrium are similar to the standard harmonic values, as shown in the main text. The SSCHA shear frequencies at SP instead change significantly. The shear mode that looks like an acoustic mode takes a value close to 2 cm-1 at $\Gamma$ (see also Fig. S1). The rest of the SSCHA phonon dispersion presents small differences between equilibrium and SP frequencies.
In addition, we performed calculations in the classical limit ($\hbar\omega/kT\rightarrow 0$) by increasing the carbon atomic masses by a factor of 600. This reduces the Debye temperature below 100 K, ensuring that all modes are thermally populated and recovering the classical Boltzmann distribution for the distorted configurations. As we can see in Fig. S5, the standard quantum SSCHA calculation and its classical limit coincide, showing quantum effects are neglegible above 100 K.
Figure S1: Unstable shear frequency at SP as a function of $1/N$, for different temperatures. The fact that it does not stabilize shows that the shear mode is present in a small region of the BZ.
Figure S2: Barrier as a function of temperature, for different supercells, just using the $N\times N\times 1$ grid to determine the harmonic part of the free energy Eq. (4). The barrier changes significantly with $N$, which means the shear mode is overrepresented. On the other hand, interpolating the harmonic part of the free energy, convergence is much better. See Fig. S3.
Figure S3: Barrier as a function of $1/N$, for different temperatures, interpolating the first term of Eq. 4. The barrier seems well converged at low temperatures and harder to converge at higher temperatures. Values are likely converged within less than 0.05 meV/atom.
Figure S4: SSCHA frequencies around the shear modes, at the equilibrium position (blue) and SP (red).
Figure S5: Comparison between the standard quantum SSCHA results and its classical limit.
S2 Shear and breathing frequency
The SSCHA auxiliary and physical frequencies as a function of temperature can be observed in Fig. S6. The frequencies are not fully converged, and the converged value might be similar to the QHA. The bulk temperature dependence (corrected with the factor of the nearest neighbor model) agrees very well with QHA.
The other mode at $\Gamma$ that involves the layers moving out of phase is the breathing mode (out-of-plane instead of in-plane). Its temperature dependence can be seen in Fig. S7. It decreases more at low temperatures compared to the shear mode and barrier. There are no measurements of the breathing mode as a function of temperature, so our calculations serve as a prediction. The SSCHA curve is very well converged, so we expect it to have a temperature dependence that is closer to the experiment relative to the QHA curve. The auxiliary frequency for $N=8$ and 10 is almost the same (Fig. S8), so the physical values are also converged (within 0.5 cm-1 at 1000 K, and less for lower temperatures).
Figure S6: Shear frequency as a function of temperature, both for the physical frequency (full lines) and the auxiliary one (dashed). The experimental value is shifted and included twice, to facilitate the comparison of the temperature dependence with the QHA and SSCHA curves. The agreement of the experimental bulk value (adjusted with the nearest neighbor model to the bilayer case) with QHA is excellent, and with SSCHA is very good as well.
Figure S7: Breathing frequency as a function of temperature using SSCHA (black circles) and QHA (red crosses). The SSCHA calculations are well converged and serve as a prediction, since there is no experimental data in the literature.
Figure S8: Convergence of the auxiliary (dashed) and physical (full) breathing frequency. The physical frequency is well-converged at $N=10$.
S3 Thermal expansion (TE)
For the SSCHA out-of-plane TE calculations, we also used the interpolation scheme with a $40\times 40$ grid. The interlayer distance and in-plane lattice parameter obtained in bulk with SSCHA can be seen in Fig. S9. The fits are used to reduce numerical error, and correspond Eq. (6) of Ref. Zorzi and Perottoni, 2021,
$b(T)=b_{0}\mathrm{exp}\left(\frac{X\theta}{e^{X/T}-1}\right),$ (S1)
which depends on parameters $b_{0}$, $X$ and $\theta$ ($b$ can be $a$ or $c_{d}$). The derivative of the fits give the coefficients of TE. In the QHA case, the TE can be calculated quickly and accurately for many temperatures, so we just use finite differences.
In Fig. S10 we compare SSCHA and QHA for bilayer graphene and bulk. The TE for both SSCHA and QHA bulk is lower relative to bilayer. In-plane, Fig. S11, we can also see a clear reduction of the TE coefficient from monolayer to bulk. In what follows, we describe in detail convergence studies we did for SSCHA, and also for QHA calculations.
Figure S9: Out-of plane (a) and in-plane (b) SSCHA thermal expansion (blue dots). The fit corresponds to Eq. (S1).
Figure S10: SSCHA and QHA out-of-plane coefficient of TE for graphite and bilayer graphene.
Figure S11: In-plane coefficient of TE using SSCHA and QHA, for monolayer, bilayer and bulk.
S3.1 Bilayer
QHA. We first considered the out-of-plane TE by fixing $a$, and then we considered the full minimization of the free energy by varying both $a$ and $c_{d}$. In Fig. S12 we can see $c_{d}(T)$ both for fixed $a$, and for the full minimization. The full minimization has a slightly larger dependence with temperature, but fixing $a$ is a good approximation. Fig. S13 is analogous for $a(T)$, with a fixed value $c_{d}=3.5$. The curves are very similar, except for a small difference at high temperatures (of about 0.0005 Å at 1000 K).
We also studied convergence with the size of the grid (Fig. S14). At $N=12$, $c_{d}(T)$ is well converged (within 0.005 Å at 1000 K), although $N\sim 20$ to achieve full convergence. The shear mode increases TE because when layers are displaced relative to each other, the interlayer distance is larger (see Fig. S15). The monotonous behavior indicates that the shear modes are overrepresented for coarser grids, and become better sampled with denser grids (having wavevectors $\mathbf{q}$ closer to $\Gamma$).
In-plane, a larger $N$ of about 50 is needed to fully achieve convergence (Fig. S16). Again, monotonous convergence suggests inadequate sampling close to $\Gamma$. Since the convergence in bulk is much faster (see below Fig. S24), we attribute the slow convergence to the flexural mode, which becomes linear (an acoustic mode) and easier to sample in bulk.
[…]
On the other hand, $a(T)$ has less variation with temperature relative to QHA and is already converged at about $N=8$ (Fig. S18).
S3.2 Bulk
QHA. In bulk, there is virtually no difference for the out-of-plane TE when fixing $a$, or considering also the minimization of $a$ (Fig. S20). In-plane (Fig. S21), there is a small difference between both fixing or varying $a$, similar to the bilayer case.
Just like in the bilayer case, relative dense grids are needed to achieve out-of-plane convergence (Fig. S22), which is again likely due to the shear mode. Regarding the extension of the supercell in the out-of-plane direction, 2 unit cells (4 layers) give good results (Fig. S23). In-plane (Fig. S24), convergence is much faster than out-of-plane.
SSCHA. Here we also used the interpolation method ($40\times 40$ grid), to go beyond the $N=8$ supercell that was possible in the direct SSCHA calculation, resulting in the blue dots of Fig. S9 (a).
In-plane, based on $N=8$ already giving converged results in bilayer (Fig. S18), and also considering that the QHA convergence in bulk (Fig. S24) is much faster than in bilayer (Fig. S16), $N=8$ should give converged SSCHA results in bulk, which corresponds to the blue dots in Fig. S9 (b).
Figure S12: Interlayer distance $c_{d}$ as a functi on of temperature in bilayer graphene using QHA, fixing the value of $a$ (dashed-blue) and doing the full minimization (full-black). In-plane TE has little effect on the out-of-plane TE.
Figure S13: Lattice parameter $a$ as a function of temperature in bilayer graphene using QHA, fixing the value of $c_{d}$ (dashed-blue) and doing the full minimization (full-black). In-plane TE has little effect on the out-of-plane TE.
Figure S14: Interlayer distance as a function of temperture in bilayer graphene, using QHA. At $N\sim 20$ $c_{d}$ is converged, although $N=12$ also gives a good result.
Figure S15: Interlayer distance profile along the bond direction, when moving a graphene layer relative to another one (analogous to Fig. 1, but for distance instead of energy). The minimum distance is at the equilibrium positions.
Figure S16: In-plane lattice parameter as a function of temperature in bilayer graphene for different $N$ using QHA. A large $N$ is needed to achieve convergence.
Figure S17: Interlayer distance as a function of temperature in bilayer graphene for different $N$ using SSCHA. Convergence is similar to the QHA case, Fig. S14, where $N=12$ gives a good value of $c_{d}(T)$.
Figure S18: In-plane lattice parameter as a function of temperature, for different $N$. At $N=8$ the curve is already well converged.
Figure S19: Interlayer distance as a function of temperature in bilayer graphene, using SSCHA, at two different values of the in-plane lattice parameter. The curves are similar to the ones of Fig. S12. In-plane TE has little effect on the out-of-plane TE.
Figure S20: Interlayer distance as a function of temperature in bulk using QHA, fixing the value of $c_{d}$ (dashed-blue) and doing the full minimization (full-black). In-plane TE has barely no effect on the out-of-plane TE.
Figure S21: In-plane lattice parameter as a function of temperature in bulk using QHA, fixing the value of $c_{d}$ (dashed-blue) and doing the full minimization (full-black). The out-of-plane TE has a slightly larger but similar effect relative to bilayer on the in-plane TE.
Figure S22: Interlayer distance as a function of temperature in graphite for different $N$ using QHA. Convergence is similar to the bilayer case.
Figure S23: Interlayer distance as a function of temperature in graphite for different $N_{z}$. $N_{z}=2$ is the value used in our calculations.
Figure S24: In-plane lattice parameter as a function of temperature in graphite using QHA, for different $N$. At $N=6$ convergence is already very good, as opposed to the bilayer case Fig. S17.

### v4

@@ III.2 Shear frequency
[…]
Figure 4: Shear frequency as a function of temperature using SSCHA (blue circles) and QHA (red crosses). There is no experimental data for the temperature dependence of the shear frequency as a function of temperature for bilayer graphene. Experimental values correspond to bulk (blue crosses) and folded 2+2 graphene (orange circles), which are corrected by the factor of a nearest neighbor model50 to compare to the bilayer case (and rigidly shifted by -5 cm-1 to facilitate the visual comparison). The agreement with the temperature dependence of the bulk values is good.
Regarding the experimental data, in Ref. 33 the measurements are done in 2+2 folded graphene (4 layers), and in Ref. 34 they are done in bulk. To compare with our bilayer calculations, we rely on a nearest layer model which has simple conversion factors between different number of layers29. Since interlayer interactions in graphite are known to be weak, this should provide a good approximation. Indeed, there are multiple references that have measured the shear frequency for different number of layers and obtained that the model describes the trend very well, like in multilayer graphene 29 and NbSe2 50. The bulk case can be readily understood: while in the bilayer case the restitutive lateral force is exerted only by the other layer, in the bulk case each layer is surrounded by two of them, so the force is doubled. Since the interatomic force constant matrix is proportional to the square of the phonon frequencies, the shear frequency in bilayer is $\sqrt{2}$ smaller than the bulk frequency in this model. Using these factors29, we obtain the adjusted experimental values of Fig. 4.
The temperature dependence of both SSCHA and QHA curves is similar to the experimental adjusted bulk values. Folded graphene values have a steeper dependence but are likely unsuitable for comparison, because layers fold with different rotational angles, while layers are aligned in standard graphite. The shear frequency has also been measured in other materials like bilayer NbSe2 in Ref. 50, or bulk $h$-BN in Ref. 51, which show a similar temperature dependence (adjusting the bulk value by a $\sqrt{2}$ factor) of roughly -0.5 cm-1/100 K.
@@ III.3 Thermal expansion
Figure 5: (a) Graphite out-of-plane coefficient of TE as a function of temperature using SSCHA and QHA. PIMD calculations39 and experimental data 37 are also included. PIMD overall agrees well, while QHA and SSCHA (with the potential of Ref. 45) underestimates the TE. (b) Graphite in-plane coefficient of TE using SSCHA and QHA. QHA again underestimates the dependence at low temperatures, and also at high temperatures. On the other hand, PIMD gives a dependence that is surprisingly very far from experimental values. SSCHA gives excellent results both at low and high temperatures, considering also the disagreement within experimental data.
To further check the adequacy of our method (the use of SSCHA with the in-plane44 and out-of-plane45 interatomic potentials), we look at the TE in the previous calculations and also in bulk graphite, and compare to experiments. The TE coefficients are given by $\alpha_{\mathrm{a}}(T)=\tfrac{1}{a}\tfrac{da}{dT}$ and $\alpha_{\mathrm{c}}(T)=\tfrac{1}{c}\tfrac{dc}{dT}$, with $a$ and $c$ the in-plane and out-of-plane lattice parameters, respectively.
In Fig. 5, we show the TE coefficient calculated with SSCHA and QHA in bulk, and compare with experimental data37, 40 and PIMD calculations 39. SSCHA results underestimate the TE , which we attribute to the interlayer potential45. This potential is fitted to energies and forces that use the many-body dispersion method in conjunction with a PBE functional, as it reproduces binding energies and interlayer distances reasonably well. However, it yields an elastic modulus along the c-axis about 15% lower than experimental values (see more details within Ref. 45). Unsurprisingly, Ref.39 obtains better results with the LCBOPII potential52 – although the TE remains somewhat below experiments37, 40 – since it uses the experimental out of plane compressibility as a fitting parameter53. The bilayer TE is also similarly underestimated (see SI). Thus, the variation of the potential around the interlayer equilibrium distance seems to be the main factor influencing the TE, as opposed to the weaker long-range interactions involving subsequent layers. It is also possible that anharmonicities, captured more accurately by PIMD relative to SSCHA, play a role at high temperatures. In any case, for the purpose of studying the barrier – the focus of our work – the problem with the LCBOPII potential is that it includes no registry effects, making the barrier essentially non-existent.
While the out-of-plane TE we obtained is not optimal, in-plane SSCHA calculations give excellent results. QHA works well at low temperatures but departs from experiments at higher temperatures. The negative TE in graphite and other layered systems is well-known and arises from flexural (out-of-plane) modes 54. We briefly summarize the mechanism here. In flexural modes, atoms are displaced out-of-plane, reducing the average projected in-plane distance, even though the 3D bond lengths remain essentially unchanged. This can also be understood intuitively via the “membrane effect”: when a layer is stretched, the out-of-plane amplitude decreases and the frequency increases, which corresponds to a negative Grüneisen parameter. See, for example, Ref. 54 for more details. We conducted a careful convergence of these results, reported in the SI, Sec. S3. PIMD calculations of Ref. 39, although they work very well out-of-plane, greatly overestimate in-plane TE, presumably due to the inadequacy of their interatomic potential. It would be interesting in future work to compare SSCHA and PIMD using the same potentials.
@@ IV Conclusions
Data availability
Additional data supporting this article has been included as part of the Supplementary Information. The SSCHA42 code used in this article is publicly available at https://sscha.eu/.
Conflicts of interest
There are no conflicts to declare.